OpenStudy (madhatter):

Which ordered pair is a solution of the equation 3x + y = –23? (5, 8) (–8, –5) (–5, –8) (8, 5)

OpenStudy (shamim):

-5,-8

OpenStudy (madhatter):

ok thanks so much i suck at this stuff

OpenStudy (anonymous):

I agree with @shamim

OpenStudy (shamim):

welcome

OpenStudy (anonymous):

-5,-8

OpenStudy (anonymous):

tn=a+(n-1)d 56=a+(8-1)5 56=a+35 a=56-35=11 An=11+(n-1)5 An=11+5n-5 An=5n+6

OpenStudy (madhatter):

??????

OpenStudy (madhatter):

@surjit